# When OpenClaw Meets Hospital: Toward an Agentic Operating System for Dynamic Clinical Workflows

### 12 Mar 2026 v1

@@ Abstract
Large language model (LLM) agents extend conventional generative models by integrating reasoning, tool invocation, and persistent memory. Recent studies suggest that such agents may significantly improve clinical workflows by automating documentation, coordinating care processes, and assisting medical decision making [1]. However, despite rapid progress, deploying autonomous agents in healthcare environments remains difficult due to reliability limitations, security risks, and insufficient long-term memory mechanisms. This work proposes an architecture that adapts LLM agents for hospital environments. The design introduces four core components: a restricted execution environment inspired by Linux multi-user systems, a document-centric interaction paradigm connecting patient and clinician agents, a page-indexed memory architecture designed for long-term clinical context management, and a curated medical skills library enabling ad-hoc composition of clinical task sequences. Rather than granting agents unrestricted system access, the architecture constrains actions through predefined skill interfaces and resource isolation. We argue that such a system forms the basis of an Agentic Operating System for Hospital, a computing layer capable of coordinating clinical workflows while maintaining safety, transparency, and auditability. This work grounds the design in OpenClaw, an open-source autonomous agent framework that structures agent capabilities as a curated library of discrete skills, and extends it with the infrastructure-level constraints required for safe clinical deployment.
1 Introduction
@@
Large language models have recently evolved from passive conversational systems into autonomous agents capable of interacting with external software environments. Agent architectures combine reasoning, planning, tool invocation, and memory to execute complex workflows that extend beyond simple text generation [2]. These capabilities have stimulated increasing interest in applying agentic systems to healthcare, where coordination across heterogeneous data sources and clinical processes is essential.
@@
Several recent studies suggest that LLM-based agents could assist with clinical tasks including diagnostic support, medical documentation, and patient communication [1]. Yet translating these systems into operational hospital environments exposes persistent structural limitations. Existing agentic frameworks assume open computing environments and routinely require broad operating-system permissions—unrestricted file-system access, external network calls, and arbitrary code execution—that are fundamentally incompatible with healthcare privacy regulations and audit requirements. Separately, dominant memory approaches based on flat vector retrieval fragment patient histories into isolated embeddings, discarding the longitudinal structure on which clinical reasoning depends and producing unreliable recall across extended care episodes.
@@
Healthcare systems impose additional constraints. Hospitals operate through structured workflows involving multiple participants, including physicians, nurses, and patients. Information exchange occurs primarily through structured documents such as clinical notes, laboratory reports, and treatment plans rather than continuous dialogue. Most current agent architectures assume a single conversational interface rather than collaborative knowledge environments. A further constraint arises from the nature of existing hospital information systems themselves. Electronic health records, clinical decision support tools, and care-coordination platforms implement care as fixed, pre-programmed workflows designed for anticipated care pathways. These systems perform well on common, well-defined protocols but fail on the long tail of clinical variability: because no two patients present identically and no disease trajectory follows a scripted sequence, any sufficiently diverse patient population will regularly surface clinical needs that no installed workflow addresses. Physicians who encounter an unusual drug interaction in a patient with a rare comorbidity profile, or patients who require guidance spanning multiple care episodes, are systematically underserved by systems that can only execute predetermined routines.
@@
Taken together, these limitations indicate that adapting agents for healthcare is not primarily a question of model capability, but of infrastructure design. We argue that agents must be embedded within a controlled computing layer specifically shaped by the safety, memory, and coordination requirements of clinical environments. This paper proposes such a layer, which we term the Agentic Operating System for Hospital, and describes its four core components: a restricted execution environment, a page-indexed memory architecture, a document-mediated multi-agent coordination model, and a medical skills library enabling ad-hoc clinical task composition.
2 Background and Motivation
@@ 2.1 LLM Agent Architectures
Recent research has explored LLM agents as modular systems integrating reasoning, tool usage, and memory. Architectures such as ReAct [3], Reflexion [4], and generative agents [5] demonstrate how language models can coordinate multi-step interactions with external environments. These frameworks typically grant agents broad access to system resources, including file systems, external APIs, and arbitrary code execution environments. While such permissiveness is acceptable in sandboxed research settings, it introduces unacceptable risks in operational environments where safety and auditability are mandatory requirements. The challenge, therefore, is not solely one of capability but of containment: how to embed agentic reasoning within a controlled execution boundary without sacrificing functional utility. Existing agent frameworks provide no principled solution to this problem, motivating the need for an infrastructure layer analogous to an operating system. One concrete example is OpenClaw, an open-source autonomous agent by Peter Steinberger that uses messaging platforms as its user interface and organises capabilities as a library of discrete skills stored as structured directories [6]. While OpenClaw demonstrates that skill-based decomposition makes agent behaviour more modular and inspectable, its design assumes a permissive, consumer-facing environment and therefore lacks the execution isolation, audit trail mechanisms, and longitudinal memory required for hospital deployment.
2.2 Retrieval-Augmented Generation and Its Limitations
@@
Several approaches have attempted to address structural limitations in RAG. MemGPT [8] introduces a hierarchical memory model inspired by OS virtual memory paging, moving context between fast in-context storage and slower external storage; however, its paging policy is designed for conversational continuity rather than document-structured clinical knowledge. GraphRAG [9] constructs a knowledge graph over a corpus and retrieves via community summaries, improving multi-hop reasoning but requiring expensive offline graph construction that does not accommodate the continuous document mutations characteristic of live patient records. RAPTOR [10] builds a recursive tree of abstractive summaries for hierarchical retrieval, but operates over static corpora and provides no mechanism for incremental index maintenance on document updates. HippoRAG [11] models retrieval as associative memory over a knowledge graph, improving cross-document reasoning, but likewise assumes a static document collection. In contrast, the page-indexed memory architecture proposed here abandons vector similarity entirely. Rather than encoding documents as embeddings, it exposes the document hierarchy to the agent through human-readable manifest files and relies on the agent’s own language-reasoning capability to navigate progressively toward relevant content. This design eliminates embedding computation overhead, requires no offline index construction, and naturally handles live document mutations without any reindexing step.
2.3 Clinical AI Deployment and Workflow Alignment
Within healthcare, substantial work has examined the requirements for integrating AI systems into clinical environments [12, 13, 14]. These studies consistently find that successful deployment requires alignment with existing institutional structures—documentation practices, professional role boundaries, and regulatory governance—rather than the open, permissive computing assumptions that underlie general-purpose agent frameworks. A distinctive feature of hospital environments is their inherently multi-participant structure: physicians, nurses, and patients each occupy defined roles and exchange information primarily through structured documents rather than continuous dialogue. This stands in sharp contrast to general-purpose agent architectures, which assume a single conversational interface between one user and one model instance. Clinical settings therefore demand a multi-agent architecture in which separate agents serve distinct participant roles and coordinate indirectly through shared document artifacts. Regulatory requirements in healthcare further demand that system behaviour be explainable and auditable, placing additional constraints on memory design, skill execution, and inter-agent communication.
@@ 2.3 Clinical AI Deployment and Workflow Alignment
2.4 Summary of Gaps and Proposed Approach
Taken together, the literature reveals four unresolved gaps. First, current agent frameworks lack a principled mechanism for restricting execution to well-defined, auditable actions. Second, existing memory systems based on flat vector retrieval are poorly suited to the hierarchical, longitudinal structure of clinical records. Third, no existing architecture explicitly models the multi-participant, document-mediated nature of hospital workflows. Fourth, existing hospital IT systems implement care as static, pre-programmed workflows that cannot adapt to the inherent long-tail variability of clinical practice; physicians and patients regularly encounter situations that fall outside installed protocols, with no mechanism for ad-hoc task execution. Our work addresses each of these gaps directly: a restricted execution environment inspired by Linux multi-user systems targets the first; a page-indexed memory architecture addresses the second; a document-centric multi-agent design addresses the third; and the agentic task-composition model—through which the agent reasons over a library of medical skills and composes them dynamically to address novel clinical requests—targets the fourth. Together these components constitute the proposed Agentic Operating System for Hospital.
3 Agentic Operating System for Hospital
@@
To support reactive coordination, each agent subscribes to a change-notification stream associated with the documents within its access scope. Internally, every document write is recorded as an append-only mutation event containing a document identifier, the modified page reference, the writing agent’s role, and a monotonically increasing version counter. A lightweight event broker dispatches these mutation events to all subscribed agents; upon receipt, an agent evaluates whether the change warrants immediate action—such as escalating an abnormal finding to a clinician—or deferred processing during the next scheduled reasoning cycle. This subscription mechanism decouples agents temporally while preserving causal ordering: because every event references a version counter, agents can reconstruct the exact sequence of document changes and detect concurrent modifications without polling. All mutation events are persisted alongside the document itself, yielding a complete, tamper-evident audit trail of every inter-agent information exchange across the system; Figure 1 summarises the overall architecture. Among the four components described above, the page-indexed memory architecture requires the most detailed treatment and is elaborated in Section 4.
4 Page-Indexed Memory Architecture
The page-indexed memory architecture organizes agent memory as a three-stage lifecycle: capture, store, and recall (Figure 2). In the capture stage, observations produced during agent interactions—sensor readings, dialogue turns, skill outputs—are written as structured entries into the document collection via the mutation mechanism described in Section 3. In the store stage, the system consolidates captured entries into a document hierarchy and maintains a manifest at each hierarchical level to summarize its contents. In the recall stage, the agent navigates this hierarchy through successive manifest reads, progressively narrowing its focus until it reaches the pages most relevant to its current reasoning task. Critically, this architecture requires no vector embeddings: navigation is driven entirely by the agent’s own language-reasoning capability applied to manifest text.
4.1 Document Hierarchy and Manifest Files
@@
4.2 Progressive Disclosure Retrieval
@@
4.3 Manifest Maintenance on Document Mutation
Each document mutation event (Section 3) may invalidate the manifest of the parent node. Upon receiving a mutation event for page under node , the system issues an incremental LLM call that rewrites only the affected child entry in , leaving all other entries unchanged. If a page is inserted or deleted, the corresponding entry is added or removed from and the description regenerated for adjacent entries if their scope changes. Manifests at ancestor nodes are updated only when the scope description of the modified subtree changes materially (e.g., a new clinical episode begins), determined by a brief LLM comparison of the old and new . This localized update policy ensures that manifest files remain accurate reflections of live document state at the cost of at most incremental LLM calls per mutation event, with no embedding recomputation at any level.
5 Clinical Workflow Integration
@@
5.1 Continuous Monitoring and Routine Follow-Up
@@
5.2 Triage and Initial Assessment
@@ 5.3 Emergency Escalation
5.4 Medication Management and Adherence Tracking
@@
5.5 Proactive Risk Identification and Preventive Intervention
@@
5.6 Cross-Specialty Coordination for Multimorbid Patients
Patients managed concurrently by multiple specialists present a coordination challenge that the document-mutation model addresses directly through shared document access and event subscription. Each specialist discipline is represented by a dedicated clinician agent—a cardiology agent, a nephrology agent, an oncology agent—each holding write access to its own specialty note documents and read access to a shared multidisciplinary care plan document. When the cardiology agent updates the antihypertensive regimen, it writes a new medication order page and appends a summary entry to the shared care plan; the resulting mutation events are delivered to all subscribed specialist agents. Each specialist agent that receives the notification performs a manifest-guided retrieval of its own relevant history—the nephrology agent navigates to recent renal function pages, the oncology agent to current chemotherapy protocol pages—and evaluates the change against its own treatment context. If a conflict is identified, such as a newly prescribed ACE inhibitor contraindicated by the patient’s current chemotherapy agent, the detecting agent appends a structured conflict alert page to the shared care plan document, triggering a further mutation event visible to all specialist agents and the coordinating clinician. The coordinating clinician agent aggregates these conflict entries, retrieves the full relevant context across all specialty manifests, and produces a reconciled care plan page that supersedes the conflicting orders. Because every write in this workflow is a versioned document mutation, the full sequence of specialist updates, conflict detections, and plan reconciliations is preserved in the audit trail with causal ordering intact, providing a transparent record equivalent to a documented multidisciplinary team meeting.
6 Safety and Governance
A key insight of the proposed architecture is that reducing the agent’s effective interface to two primitive operations—file read and file write—allows the entire safety and governance problem to be delegated to the security machinery of the host operating system rather than implemented in bespoke application-level logic.
6.1 Minimal Trusted Computing Base

### 21 Mar 2026 v2

@@ Abstract
Large language model (LLM) agents extend generative models with reasoning, tool use, and persistent memory, thereby enabling the automation of complex tasks. In healthcare, such systems could support documentation, care coordination, and clinical decision making. Their reliable deployment in hospitals, however, remains constrained by safety risks, limited transparency, and inadequate mechanisms for handling longitudinal clinical context. Here we propose an architecture that adapts LLM agents to hospital environments. The design comprises four components: a restricted execution environment inspired by multi-user operating systems, a document-centric interaction model linking patient and clinician agents, a page-indexed memory architecture for longitudinal context management, and a curated library of composable medical skills. Implemented on top of OpenClaw, an open-source agent orchestration framework, this design provides the basis for an Agentic Operating System for Hospitals: a computing layer for coordinating clinical workflows while preserving safety, transparency, and auditability. To evaluate the memory component, we introduce manifest-guided retrieval for hierarchical navigation of longitudinal patient records. In a benchmark derived from the MIMIC-IV dataset (v2.2) comprising 100 de-identified patient records and 300 clinical queries stratified across three difficulty tiers (100 per tier), manifest-guided retrieval matched a metadata-filtered RAG baseline on overall recall (0.877 versus 0.876) while achieving 2.2 higher precision (0.779 versus 0.352) and retrieving fewer documents; on tier-3 longitudinal queries, manifest recall was 21% higher (0.846 versus 0.701), confirming that LLM-guided hierarchical navigation is most valuable when queries span multiple care episodes. These results outline a practical path toward hospital-scale agentic infrastructure.
keywords:
large language models; clinical workflow automation; manifest-guided retrieval; healthcare AI; agentic operating system1 Introduction
@@
Recent work has extended large language models (LLMs) into autonomous agents capable of interacting with external software environments. Agent architectures combine reasoning, planning, tool use, and memory to execute workflows that extend beyond text generation alone [1]. These capabilities have stimulated growing interest in healthcare, where coordination across heterogeneous data sources, clinical documents, and care processes is central to everyday practice.
@@
Several recent studies suggest that LLM-based agents could assist with diagnostic support, clinical documentation, and patient communication [2]. Translating such systems into operational hospital settings, however, exposes persistent structural limitations. Existing agentic frameworks generally assume open computing environments and broad operating-system permissions, including unrestricted file-system access, external network calls, and arbitrary code execution. These assumptions are fundamentally misaligned with healthcare privacy regulation and audit requirements. At the same time, dominant memory approaches based on flat vector retrieval fragment patient histories into isolated embeddings, discarding the longitudinal structure on which clinical reasoning depends and weakening recall across extended care episodes.
@@
Hospital environments impose additional constraints. Care is delivered through structured workflows involving multiple participants, including physicians, nurses, and patients. Information exchange occurs primarily through structured artifacts such as clinical notes, laboratory reports, and treatment plans rather than through continuous dialogue. Most current agent architectures, by contrast, assume a single conversational interface rather than a collaborative document-centered environment. A further limitation arises from the design of existing hospital information systems. Electronic health records, clinical decision support tools, and care-coordination platforms typically encode care as fixed, pre-programmed workflows designed for anticipated pathways. These systems perform well for common, well-defined protocols but are poorly suited to the long tail of clinical variability, because patient presentations and disease trajectories rarely conform to predefined routines. Physicians confronting unusual comorbidity profiles, or patients seeking guidance that spans multiple episodes of care, are therefore often left beyond the reach of installed workflows.
@@
Taken together, these observations suggest that adapting agents for healthcare is not primarily a problem of model capability, but of infrastructure design. We argue that agents should operate within a controlled computing layer purpose-built to satisfy the safety, memory, and coordination requirements of clinical environments. We therefore introduce the concept of an Agentic Operating System for Hospitals, a computing layer designed to support safe agent deployment in clinical settings. We describe four core components addressing execution control, longitudinal memory organization, document-mediated coordination, and composable clinical skills.
2 Related Work
@@ 2.1 LLM agent architectures
Recent research has explored LLM agents as modular systems integrating reasoning, tool use, and memory. Architectures such as ReAct [3], Reflexion [4], and generative agents [5] demonstrate how language models can coordinate multi-step interactions with external environments. These frameworks typically grant agents broad access to system resources, including file systems, external APIs, and arbitrary code execution environments. While such permissiveness is acceptable in sandboxed research settings, it introduces unacceptable risks in operational environments where safety and auditability are mandatory requirements. The challenge, therefore, is not solely one of capability but of containment: how to embed agentic reasoning within a controlled execution boundary without sacrificing functional utility. Existing agent frameworks provide no principled solution to this problem, motivating the need for an infrastructure layer analogous to an operating system. One concrete example is OpenClaw, an open-source agent orchestration framework designed for complex, multi-step tasks [6]. Unlike tools-and-skills integration that many agent frameworks already support, OpenClaw’s distinguishing contributions lie in its persistent memory architecture and its heartbeat mechanism: a scheduled task system that enables agents to autonomously initiate actions at defined intervals without requiring an explicit user trigger. Together, these two mechanisms allow agents to maintain coherent long-term context across sessions and to proactively monitor or act upon evolving conditions—capabilities that are essential for any system expected to operate continuously in a dynamic environment. However, OpenClaw’s design targets a permissive, consumer-facing context and therefore lacks the execution isolation, audit trail mechanisms, and clinically structured longitudinal memory required for hospital deployment.
2.2 Retrieval-augmented generation in clinical contexts
@@
Several approaches have attempted to address structural limitations in RAG. MemGPT [8] introduces a hierarchical memory model inspired by OS virtual memory paging, moving context between fast in-context storage and slower external storage; however, its paging policy is designed for conversational continuity rather than document-structured clinical knowledge. GraphRAG [9] constructs a knowledge graph over a corpus and retrieves via community summaries, improving multi-hop reasoning but requiring expensive offline graph construction that does not accommodate the continuous document mutations characteristic of live patient records. RAPTOR [10] builds a recursive tree of abstractive summaries for hierarchical retrieval, but operates over static corpora and provides no mechanism for incremental index maintenance on document updates. HippoRAG [11] models retrieval as associative memory over a knowledge graph, improving cross-document reasoning, but likewise assumes a static document collection. By contrast, we propose a page-indexed memory architecture that abandons vector similarity entirely. Rather than encoding documents as embeddings, it exposes the document hierarchy to the agent through human-readable manifest files and relies on the agent’s own language-reasoning capability to navigate progressively toward relevant content. This design eliminates embedding computation overhead, requires no offline index construction, and naturally handles live document mutations without any reindexing step.
2.3 Clinical workflow integration and deployment challenges
Within healthcare, substantial work has examined the requirements for integrating AI systems into clinical environments [12, 13, 14]. These studies consistently show that successful deployment depends on alignment with existing institutional structures—documentation practices, professional role boundaries, and regulatory governance—rather than with the open, permissive computing assumptions that underlie general-purpose agent frameworks. A distinctive feature of hospital environments is their inherently multi-participant structure: physicians, nurses, and patients each occupy defined roles and exchange information primarily through structured documents rather than continuous dialogue. This stands in sharp contrast to general-purpose agent architectures, which assume a single conversational interface between one user and one model instance. Clinical settings therefore demand a multi-agent architecture in which separate agents serve distinct participant roles and coordinate indirectly through shared document artifacts. Regulatory requirements in healthcare further demand that system behaviour be explainable and auditable, placing additional constraints on memory design, skill execution, and inter-agent communication.
@@ 2.3 Clinical workflow integration and deployment challenges
2.4 Research gaps
Taken together, the literature reveals four unresolved gaps. First, current agent frameworks lack a principled mechanism for restricting execution to well-defined, auditable actions. Second, existing memory systems based on flat vector retrieval are poorly suited to the hierarchical, longitudinal structure of clinical records. Third, no existing architecture explicitly models the multi-participant, document-mediated nature of hospital workflows. Fourth, existing hospital IT systems implement care as static, pre-programmed workflows that cannot adapt to the inherent long-tail variability of clinical practice; physicians and patients regularly encounter situations that fall outside installed protocols, with no mechanism for ad-hoc task execution. Our work addresses each of these gaps directly: a restricted execution environment inspired by Linux multi-user systems targets the first; a page-indexed memory architecture addresses the second; a document-centric multi-agent design addresses the third; and the agentic task-composition model—through which the agent reasons over a library of medical skills and composes them dynamically to address novel clinical requests—targets the fourth. Together these components constitute the proposed Agentic Operating System for Hospitals.
3 System Architecture
3.1 Restricted execution and skills coordination
@@
To support reactive coordination, each agent subscribes to a change-notification stream associated with the documents within its access scope. Internally, every document write is recorded as an append-only mutation event containing a document identifier, the modified page reference, the writing agent’s role, and a monotonically increasing version counter. A lightweight event broker dispatches these mutation events to all subscribed agents; upon receipt, an agent evaluates whether the change warrants immediate action—such as escalating an abnormal finding to a clinician—or deferred processing during the next scheduled reasoning cycle. This subscription mechanism decouples agents temporally while preserving causal ordering: because every event references a version counter, agents can reconstruct the exact sequence of document changes and detect concurrent modifications without polling. To prevent logically inconsistent concurrent updates, write-capable reasoning cycles are serialized at the patient scope through a short-lived filesystem lease. Before an agent begins a read–reason–write operation over a patient’s document subtree, it must first acquire that patient’s lock directory using an atomic filesystem primitive such as mkdir or open(O_CREAT|O_EXCL). The lease record stores the agent identity, acquisition time, and last-heartbeat timestamp; the agent renews the lease while running and releases it immediately after the final page append completes. If the agent crashes, stalls, or exceeds a bounded execution window, the lease expires automatically and can be safely reclaimed after a timeout. This policy intentionally chooses coarse-grained locking: all writes affecting a single patient are serialized for a short interval, while unrelated patients remain fully parallelizable across the system. Because the mechanism is implemented directly through standard Linux filesystem semantics, it requires no distributed lock service and remains fully auditable through the same OS-native monitoring stack as ordinary document writes. All mutation events are persisted alongside the document itself, yielding a complete, tamper-evident audit trail of every inter-agent information exchange across the system; Figure 1 summarises the overall architecture. Among the four components described above, the page-indexed memory architecture requires the most detailed treatment and is elaborated in the following subsection.
3.2 Page-indexed memory architecture
The page-indexed memory architecture organizes agent memory as a three-stage lifecycle: capture, store, and recall (Figure 2). In the capture stage, observations produced during agent interactions—sensor readings, dialogue turns, skill outputs—are written as structured entries into the document collection via the mutation mechanism described above. In the store stage, the system consolidates captured entries into a document hierarchy and maintains a manifest at each hierarchical level to summarize its contents. In the recall stage, the agent navigates this hierarchy through successive manifest reads, progressively narrowing its focus until it reaches the pages most relevant to its current reasoning task. Critically, this architecture requires no vector embeddings: navigation is driven entirely by the agent’s own language-reasoning capability applied to manifest text.
Document hierarchy and manifest files.
@@ 3.2 Page-indexed memory architecture
Progressive disclosure retrieval.
@@
Manifest maintenance on document mutation.
Each document mutation event (described above) may invalidate the manifest of the parent node. Upon receiving a mutation event for page under node , the system issues an incremental LLM call that rewrites only the affected child entry in , leaving all other entries unchanged. If a page is inserted or deleted, the corresponding entry is added or removed from and the description regenerated for adjacent entries if their scope changes. Manifests at ancestor nodes are updated only when the scope description of the modified subtree changes materially (e.g., a new clinical episode begins), determined by a brief LLM comparison of the old and new . This localized update policy ensures that manifest files remain accurate reflections of live document state at the cost of at most incremental LLM calls per mutation event, with no embedding recomputation at any level.
3.3 Clinical workflow integration
@@
Continuous monitoring and routine follow-up.
@@
Triage and initial assessment.
@@
Medication management and adherence tracking.
@@
Proactive risk identification and preventive intervention.
@@
Cross-specialty coordination for multimorbid patients.
Patients managed concurrently by multiple specialists present a coordination challenge that the document-mutation model addresses directly through shared document access and event subscription. Each specialist discipline is represented by a dedicated clinician agent—for example, a cardiology agent, a nephrology agent, or an oncology agent—each holding write access to its own specialty-note documents and read access to a shared multidisciplinary care plan document. When the cardiology agent updates the antihypertensive regimen, it writes a new medication order page and appends a summary entry to the shared care plan; the resulting mutation events are delivered to all subscribed specialist agents. Each specialist agent that receives the notification performs a manifest-guided retrieval of its own relevant history—the nephrology agent navigates to recent renal function pages, the oncology agent to current chemotherapy protocol pages—and evaluates the change against its own treatment context. If a conflict is identified, such as a newly prescribed ACE inhibitor contraindicated by the patient’s current chemotherapy agent, the detecting agent appends a structured conflict alert page to the shared care plan document, triggering a further mutation event visible to all specialist agents and the coordinating clinician. The coordinating clinician agent aggregates these conflict entries, retrieves the full relevant context across all specialty manifests, and produces a reconciled care plan page that supersedes the conflicting orders. Because every write in this workflow is a versioned document mutation, the full sequence of specialist updates, conflict detections, and plan reconciliations is preserved in the audit trail with causal ordering intact, providing a transparent record equivalent to a documented multidisciplinary team meeting.
3.4 Safety and governance
A key insight of the proposed architecture is that reducing the agent’s effective interface to two primitive operations—file read and file write—allows most of the safety and governance problem to be delegated to the security machinery of the host operating system rather than implemented as bespoke application logic.
Minimal trusted computing base.